# AudioBench: A Universal Benchmark for Audio Large Language Models

### v2

@@ Abstract
We introduce AudioBench, a new benchmark designed to evaluate audio large language models (AudioLLMs). AudioBench encompasses 8 distinct tasks and 26 carefully selected or newly curated datasets, focusing on speech understanding, voice interpretation, and audio scene understanding. Despite the rapid advancement of large language models, including multimodal versions, a significant gap exists in comprehensive benchmarks for thoroughly evaluating their capabilities. AudioBench addresses this gap by providing relevant datasets and evaluation metrics. In our study, we evaluated the capabilities of four models across various aspects and found that no single model excels consistently across all tasks. We outline the research outlook for AudioLLMs and anticipate that our open-source code, data, and leaderboard will offer a robust testbed for future model developments. [^1] https://github.com/AudioLLMs/AudioBench
@@ 1 Introduction
Although instruction-following audio-language models (or speech-text models) have attracted significant interest from both industry and research communities, their evaluation datasets are notably different. For instance, the Qwen-Audio Chu et al. (2023) model was tested on 12 distinct datasets, whereas the SALMONN Tang et al. (2024) model used 15 different datasets, with only two datasets common between them. Although WavLLM Hu et al. (2024a) has been compared with earlier models, the scope of evaluation tasks remains limited. Another limitation arises from the evaluation set; similar to text-based LLMs Chiang et al. (2024), evaluations predominantly rely on existing datasets and metrics Hendrycks et al. (2021). Given that these models follow instructions, traditional evaluation metrics can not fully meet the requirements and could introduce bias in evaluation. Consequently, it is crucial to develop innovative benchmarks for AudioLLMs. Since their ultimate goal is to handle diverse audio inputs and respond accurately to user queries, the evaluation set should include not only traditional speech tasks but also new datasets and metrics that better reflect real-world applications and assess the models’ ability to follow instructions.
Figure 1: Overview of the datasets in AudioBench.
@@
In this study, we introduce AudioBench as a comprehensive evaluation benchmark specifically designed for general instruction-following audio-language models, which we refer to as AudioLLMs. These models are designed to take comprehension of general audio inputs. These models are designed to comprehend general audio inputs. The tasks should include, but are not limited to, understanding speech semantics, interpreting environmental sounds, and recognizing paralinguistic features, as well as combinations of these elements. Therefore, as shown in Figure 1, we incorporate 8 tasks and 26 datasets covering these aspects, with 7 of the datasets being newly adapted or collected to address gaps in the existing datasets.
@@
For a fair comparison, our benchmark incorporates a variety of prompt templates to evaluate model compatibility with different instructions. Additionally, we vary the input lengths from seconds to minutes to better assess model performance on longer audio sequences. Our test suite includes 400+ hours of audio, distributed across 100k+ samples. The AudioBench toolkit will be made available, enabling the evaluation of future models simply by implementing a customized model inference program.
@@
We also explore evaluation metrics for the open-ended generation of AudioLLMs, which primarily operate as prompt-based, instruction-following models that integrate user queries into their output generation, conditioned on audio inputs. Unlike text-based large language models (LLMs), which can be accessed with multiple-choice questions, AudioLLMs are designed to manage free-format queries. This introduces difficulties in quantitatively assessing their outputs, as using LLMs like GPT-4 as judges can be costly and subject to instability due to version updates. Therefore, we investigate open-source model-as-judge methods and examine their correlation with GPT-4 models. Our findings indicate that LLaMA-3 demonstrates a higher correlation compared to Prometheus-2 Kim et al. (2024) which is specifically fine-tuned for evaluation purposes.
@@
We conducted evaluations on four open-sourced models to offer a comprehensive review of various solutions, including both end-to-end and cascaded models. Our benchmark results indicate that none of the models performed exceptionally well across all criteria, highlighting significant opportunities for future advancements of AudioLLMs with our benchmark. To guide these developments, we also include a detailed discussion on potential future directions for AudioLLMs.
@@ 2 Related Work
Multimodal LLM Benchmark. Recent progress in both text-based and multimodal LLMs has spurred the creation of a series of benchmarks designed to monitor and direct their development efficiently. Our research is consistent with earlier benchmarks on multimodal LLMs, primarily concentrating on visual components. There is a lack of comprehensive benchmarks for multimodal LLMs that provide a holistic understanding of audio inputs including speech, ambient sounds, paralinguistic features, etc. When classifying the models targeted for evaluation, we can categorize them into several distinct types: Text-based LLMs, Vision LLMs, and Audio LLMs.
@@
The evaluation of Vision-LLMs is multifaceted, encompassing aspects such as multidisciplinary understanding Yue et al. (2023); Zhang et al. (2024b); Hu et al. (2024b), perception tests Yu et al. (2023); Liu et al. (2023), mathematical reasoning Li et al. (2024a); Zhang et al. (2024a), and video understanding Li et al. (2023); Ning et al. (2023); Liu et al. (2024); Fu et al. (2024). Notably, video understanding benchmarks assess modalities that include both visual and audio elements. However, these benchmarks cannot be directly applied to audio understanding, as they predominantly rely on visual inputs with audio serving merely as a supplementary feature. Consequently, there is an urgent need to standardize benchmarks for Audio LLMs that specifically measure model performance with audio context. Such benchmarks are essential for advancing research and improving the capabilities of models that focus on audio processing and understanding.
@@
Advancements in AudioLLMs. Although no comprehensive evaluation benchmarks currently exist, we observe various multimodal foundation models that enhance speech and audio understanding, utilizing either cascaded methods or integrated multitask optimization. Some models specialize in Automatic Speech Recognition (ASR) Shu et al. (2023); Ma et al. (2024) or other non-speech audio tasks Kong et al. (2024). However, ideal speech-text foundation models often employ multitasking approaches. Notable examples include AudioGPT Huang et al. (2023), SpeechGPT Zhang et al. (2023a), SALMONN Tang et al. (2024), Qwen-Audio Chu et al. (2023), AudioPALM Rubenstein et al. (2023), LTU-AS Gong et al. (2023), VioLA Wang et al. (2023), LLaSM Shu et al. (2023), Pengi Deshmukh et al. (2023), WavLLM Hu et al. (2024a), UniverSLU Arora et al. (2024), and SpeechVerse Das et al. (2024), each focusing on a subset of tasks tailored to their specific use cases. In this study, we focus on three models—WavLLM, Qwen-Audio, and SALMONN—that represent a broader range of task-handling capabilities.
3 AudioBench
3.1 Main Concept
@@
The design of AudioLLMs is to interpret audio content and respond flexibly to user queries, handling tasks from basic speech recognition to complex reasoning. To evaluate these models effectively, we should assess their ability to understand not only speech but also environmental sounds and paralinguistic features such as emotion, gender, and accent. Therefore, our evaluation involves designing benchmarks that cover three key aspects: 1) speech understanding, focusing on the semantic content of audio and including tasks such as Automatic Speech Recognition (ASR), Speech Question Answering (SQA), and Speech Instruction (SI); 2) audio scene understanding, focusing on non-human audio sounds and including tasks such as Audio Question Answering (AQA) and Audio Captioning (AC); and 3) voice understanding, focusing on non-speech but human-related information and including paralinguistic tasks such as Emotion Recognition (ER), Accent Recognition (AR), and Gender Recognition (GR). For each task, we compile existing datasets or develop new ones tailored to the requirements for comprehensive understanding and reasoning. Figure 1 provides an overview of the tasks and the datasets used in the initial setup.
@@
In the current benchmark, our focus is on English audio understanding, where most models have not yet demonstrated satisfactory results. Meanwhile, multilingual understanding, which includes speech translation and code-switching, represents a useful application that could be integrated into the benchmarking datasets as part of an ongoing development effort.
@@
Category Dataset Name # of Samples Total Hours Avg Len (s) Min Len (s) Max Len (s) Speech Understanding ASR LibriSpeech-Clean 2,617 5.40 7.43 1.28 34.96 LibriSpeech-Other 2,935 5.34 6.55 1.47 34.51 CommonVoice-15 16,348 26.95 5.93 1.34 105.67 PeoplesSpeech 32,603 59.20 6.54 1.00 99.91 GigaSpeech 18,650 35.09 6.77 1.00 22.03 Tedlium3 1,142 2.61 8.24 1.07 32.55 Tedlium3-Longform 11 2.61 856.51 340.96 1,596.73 Earning-21 44 39.26 3,212.41 1,097.10 5,740.61 Earning-22 125 119.88 3,452.71 874.73 7,407.05 SQA CN-College-Listen 2,271 13.3 21.09 5.76 137.82 SLUE-P2-SQA5 408 4.5 39.85 13.00 40.0 DREAM-TTS 1,913 18.1 34.14 3.14 261.85 Public-SG-SpeechQA 688 7.6 39.86 15.78 95.71 SI OpenHermes-Audio 100 0.16 5.95 2.04 15.77 ALPACA-Audio 100 0.12 4.32 1.80 8.85 Audio Scene Understanding AQA Clotho-AQA 2,251 14.1 22.59 15.03 29.97 WavCaps-QA 304 0.87 10.28 1.0 30.62 AudioCaps-QA 313 0.86 9.86 3.27 10.00 AC WavCaps 1,730 4.9 10.22 1.00 30.97 AudioCaps 4,400 12.1 9.86 1.74 10.0 Voice Understanding ER IEMOCAP-Emotion 1,002 1.3 4.51 0.75 24.12 MELD-Sentiment 2,610 2.4 3.35 0.13 304.96 MELD-Emotion 2,610 2.4 3.35 0.13 304.96 AR VoxCeleb1-Accent 4,874 11.2 8.27 3.96 69.04 GR VoxCeleb1-Gender 4,874 11.2 8.27 3.96 69.04 IEMOCAP-Gender 1,002 1.26 4.55 0.69 26.77
Table 1: Datasets from the AudioBench are entirely in English and maintain a uniform sampling rate of 16 kHz. Datasets highlighted in gray and green represent newly expanded datasets and those introduced specifically for this study, respectively. Full references can be found in Figure 4.
@@ 3.2 Evaluation Setup
From a use case perspective, we expect LLMs to respond to user queries and generate natural responses. Therefore, we evaluate the models in as many free-generation scenarios as possible. For instance, a pair ($Audio$, $Query$) is presented to the model to elicit a ($Model\_Answer$), which is then compared with a reference answer for relevant tasks. This approach contrasts with traditional speech benchmarks like SUPERB Yang et al. (2024) and Dynamic-SUPERB Huang et al. (2024a), where a supervised dataset is specifically fine-tuned for a downstream task and evaluated in a single-task setting.
Therefore, we employed the model-as-judge approach for all tasks except for ASR, where word error rate (WER) is used as the sole metric, and audio captioning, where the METEOR score is utilized as an additional measure. In subsequent experiments, we explored the impact and best practices of the model-as-judge method and validated a reasonable solution.
Moreover, we found that models can be less robust to diverse instruction. This phenomenon also applies to text-based LLMs Zhu et al. (2023); Wang et al. (2024) and we found it even more severe for AudioLLMs as additional modality presents more challenges in the fusion process. Therefore, we propose to leverage multiple instructions to evaluate a single dataset if the query is not diverse enough. This applies to tasks like ASR, ER, AR, GR, and AC. We present this analysis in Section 4.
Figure 2: Sensitivity tests were conducted on three different prompts, revealing that SALMONN is the least robust when faced with varying prompt templates for the same task.
@@ 3.3.1 Speech Understanding
Figure 3: Correlation results for three benchmark datasets across three variations of the model-as-judge approach were analyzed using Spearman’s rank correlation. The findings show that Llama-3-70B-Instruct has the highest correlation with GPT-4.
Speech Question Answering (SQA). The task involves responding to questions based on speech-related audio content. However, there is a lack of suitable datasets. In this part, we involve both monologue understanding and dialogue understanding tasks and curated 3 new datasets. Our collection includes both synthesized and real speech Examples can be found in Table 5.
First, we collected questions from the English listening comprehension section of China’s national college entrance examinations, aimed at assessing students’ ability to understand English in both academic and everyday situations. We manually compiled 271 questions and added them to 2000 questions sourced from Hu et al. (2024a), which includes questions from approximately 130 different exam sets. In line with the free-text QA format, we presented the questions without multiple-choice options. The dataset is named as CN-College-Listen.
DREAM-TTS dataset is built upon text-based dialogue comprehension datasets DREAM Sun et al. (2019). We leverage one SOTA TTS engine - XTTS__\__v2 Casanova et al. (2024) to convert the text input into spoken format. The genders are keep consistent with the 60 diverse speakers randomly selected.
For the Public-SG-SpeechQA dataset, we chose four public speaking videos from Singaporean politicians, accompanied by clean transcriptions. These transcriptions were manually segmented, and five questions per segment were generated using LLMs. Each question and its corresponding reference answer were manually reviewed to ensure their validity, where 30% of the samples are discarded. Ultimately, this process resulted in the collection of 688 speech question-answer pairs for evaluation.
Speech Instruction (SI). This task evaluates whether the model can directly follow instructions provided through speech input, mirroring a natural human-computer interaction. Specifically, the question is delivered via audio, and the model is expected to understand and generate responses in the appropriate text format. This approach offers unique benefits. For instance, by incorporating paralinguistic information such as emotions, the model can adapt its responses based on the user’s emotional state. However, finding suitable testing examples has been a challenge. In this study, we use a TTS system Casanova et al. (2024) to convert an existing instruction-following dataset into the spoken format. We then involve humans in selecting instances where the speech is accurately synthesized and the content is suitable as spoken instructions. As a result, 10.5% samples are kept and the examples can be found in Table 8.
@@
Dataset Name Metrics AudioLLMs Whisper+Llama3 SALMONN Qwen-Audio WavLLM Speech Understanding LibriSpeech-Clean WER 45.28 2.20 2.05 1.83 LibriSpeech-Other 33.76 4.40 4.88 3.71 CommonVoice-15 44.44 11.30 14.84 9.89 PeoplesSpeech 35.05 30.59 37.51 14.55 GigaSpeech 14.42 13.18 16.02 9.52 Tedlium3 13.21 4.28 5.43 3.81 Tedlium3-Longform 25.88 34.16 46.51 4.94 Earning-21 31.81 39.39 65.64 11.77 Earning-22 35.87 54.37 67.22 15.60 CN-College-Listen M.J.70B 1.95 2.15 2.51 2.96 SLUE-P2-SQA5 3.95 3.84 4.16 4.19 DREAM-TTS 1.79 1.91 2.50 2.82 Public-SG-SpeechQA 2.77 2.97 2.98 3.25 OpenHermes-Audio 1.05 1.14 1.07 2.82 ALPACA-Audio 0.85 1.12 1.15 3.37 Audio Scene Understanding Clotho-AQA M.J.70B 2.75 2.88 2.16 1.40 WavCaps-QA 2.31 1.97 1.34 0.83 AudioCaps-QA 2.39 2.41 1.34 0.77 WavCaps 1.25 1.41 0.49 0.23 AudioCaps 1.83 2.11 0.50 0.24 WavCaps METEOR 21.36 21.75 9.78 14.52 AudioCaps 22.79 22.91 6.70 7.15 Voice Understanding IEMOCAP-Emotion M.J.70B 2.33 1.60 2.43 1.85 MELD-Sentiment 1.82 1.72 1.65 2.26 MELD-Emotion 1.60 1.46 1.45 1.88 VoxCeleb1-Accent 1.50 2.25 1.87 1.19 VoxCeleb1-Gender 3.98 3.59 2.96 2.87 IEMOCAP-Gender 2.02 2.55 2.01 2.51
Table 2: Main evaluation results of various AudioLLMs and one cascade model. M.J.70B refers to Model-as-Judge with Llama-3-70B-Instruct model (Score is from 0-lowest to 5-highest). METEOR scores are common metrics for audio captioning Banerjee and Lavie (2005). WER is the word error rate for ASR tasks. When outperforming end-to-end models, the Whisper+Llama3 model is marked with underline.

### v3

@@ Abstract
We introduce AudioBench, a universal benchmark designed to evaluate Audio Large Language Models (AudioLLMs). It encompasses 8 distinct tasks and 26 datasets, among which, 7 are newly proposed datasets. The evaluation targets three main aspects: speech understanding, audio scene understanding, and voice understanding (paralinguistic). Despite recent advancements, there lacks a comprehensive benchmark for AudioLLMs on instruction following capabilities conditioned on audio signals. AudioBench addresses this gap by setting up datasets as well as desired evaluation metrics. Besides, we also evaluated the capabilities of five popular models and found that no single model excels consistently across all tasks. We outline the research outlook for AudioLLMs and anticipate that our open-sourced evaluation toolkit, data, and leaderboard will offer a robust testbed for future model developments. [^1] Our Toolkit and Leaderboard can be found here.
@@ 1 Introduction
Although instruction-following audio-language models have attracted significant interest from both industry and research communities, their reported evaluation datasets are notably different. For instance, the Qwen-Audio-Chat Chu et al. (2023) model has tested on 12 datasets, whereas the SALMONN Tang et al. (2024) model used 15 datasets, with only two datasets in common. Although WavLLM Hu et al. (2024a) has been compared with some earlier models, the scope of evaluation tasks remains limited, which makes the comparison less justified. Another limitation arises from the evaluation set; similar to text-based LLMs Chiang et al. (2024), evaluations predominantly rely on previous datasets and metrics Hendrycks et al. (2021). Given that these models are expected to respond flexibly following the instructions, conventional evaluation metrics are not well prepared for this purpose. Consequently, it is crucial to develop innovative benchmarks for AudioLLMs. Since their ultimate goal is to handle diverse audio inputs and respond accurately to user queries, the evaluation set should include not only traditional speech tasks but also new datasets and metrics that better reflect their particular nature and assess the instruction-following capability of models, such as zero-shot speech question answering, etc.
Figure 1: Overview of AudioBench datasets.
@@
In this study, we introduce AudioBench as a comprehensive evaluation benchmark specifically designed for general instruction-following Audio Large Language Models, which we refer to as AudioLLMs. These models are designed to comprehend general audio inputs and user queries. The tasks should include but are not limited to, understanding speech semantics, interpreting environmental sounds, and recognizing paralinguistic features, as well as combinations of these elements. Therefore, as shown in Figure 1, we incorporate 8 tasks and 26 datasets covering these aspects, with 7 of the datasets being newly adapted or collected to address gaps in the existing datasets.
@@
For a more comprehensive assessment of the models, our benchmark incorporates a variety of prompt templates to evaluate model compatibility with different instructions. This is because we found that some models are more robust to seen instruction while do not generalize to unseen ones which could lead to robustness issues for deployment. Additionally, we vary the input lengths from seconds to minutes to better assess model performance on longer audio sequences. Our test suite includes 400+ hours of audio, distributed across 100k+ samples. Our AudioBench toolkit is available, enabling the evaluation of future AudioLLMs by simply implementing a customized model inference program.
@@
We also explored and decided what evaluation metrics to leverage, particularly for open-ended generation of AudioLLMs. Inspired by text-based and vision-based large language models (LLMs), which can be accessed with multiple-choice and open-ended questions, AudioLLMs are designed to manage diverse instruction and need specific metrics to cover the unconstrained generation space. This introduces difficulties in quantitatively assessing their outputs. Most practices go to model-as-judge methods Padlewski et al. (2024); Yang et al. (2024a). However, cloud-sourced models, e.g. GPT-4, can be costly and subject to instability introduced by commercial version updates. Therefore, we investigate open-source model-as-judge methods and validate the best practices by correlation studies. Our findings indicate that LLaMA-3-70B-Instruct model demonstrates a higher correlation compared to judgment models like Prometheus-2 Kim et al. (2024) which is specifically fine-tuned for evaluation purposes.
@@
Finally, we evaluated 5 models to offer a comprehensive review of various solutions, including both end-to-end and cascaded models. Our benchmark results indicate that none of the models performed exceptionally well across all criteria, highlighting significant opportunities for future advancements in model capabilities. To guide these developments, we also include an in-depth discussion on future challenges and directions.
@@ 2 Related Work
Multimodal LLM Benchmark. Recent progress in both text-based and multimodal LLMs has spurred the creation of a series of benchmarks designed to monitor and direct their development efficiently. Our research is consistent with earlier benchmarks on multimodal LLMs, primarily concentrating on visual components. There is a lack of comprehensive benchmarks for multimodal LLMs that provide a holistic understanding of audio inputs including speech, ambient sounds, paralinguistic features, etc. When classifying the models targeted for evaluation, we can categorize them into several distinct types: Text-LLMs, Vision-LLMs, and Audio-LLMs.
@@
The evaluation of Vision-LLMs is multifaceted, encompassing aspects such as multidisciplinary understanding Yue et al. (2023); Zhang et al. (2024b); Hu et al. (2024b), perception tests Yu et al. (2023); Liu et al. (2023), mathematical reasoning Li et al. (2024a); Zhang et al. (2024a), and video understanding Li et al. (2023); Ning et al. (2023); Liu et al. (2024); Fu et al. (2024). Notably, video understanding benchmarks assess modalities including both visual and audio elements. However, these benchmarks cannot be directly applied to audio understanding, as they predominantly rely on visual inputs with audio serving merely as a supplementary feature. Consequently, there is an urgent need to standardize benchmarks for AudioLLMs that specifically measure model performance with audio context. Such benchmarks are essential for advancing research and improving the capabilities of models that focus on audio processing and understanding.
@@
Advancements in AudioLLMs. Although no comprehensive evaluation benchmarks currently exist, we observe various multimodal foundation models that enhance speech and audio understanding, utilizing either cascaded methods or integrated multitask optimization. Some models specialize in Automatic Speech Recognition (ASR) Shu et al. (2023); Ma et al. (2024) or other non-speech audio tasks Kong et al. (2024). However, ideal speech-text foundation models often employ multitasking approaches. Notable examples include AudioGPT Huang et al. (2023), SpeechGPT Zhang et al. (2023a), SALMONN Tang et al. (2024), Qwen-Audio Chu et al. (2023), AudioPALM Rubenstein et al. (2023), LTU-AS Gong et al. (2023), VioLA Wang et al. (2023), LLaSM Shu et al. (2023), Pengi Deshmukh et al. (2023), WavLLM Hu et al. (2024a), UniverSLU Arora et al. (2024), SpeechVerse Das et al. (2024), and Qwen2-Audio Chu et al. (2024) each focusing on a subset of tasks tailored to their specific use cases. In this study, we focus on four models—WavLLM, Qwen-Audio-Chat, SALMONN, and Qwen2-Audio-Instruct—that represent a broader range of task-handling capabilities.
3 AudioLLM Benchmark (AudioBench)
3.1 Core Concept
@@
The design of AudioLLMs is to interpret audio content and respond flexibly to user queries, handling tasks from basic speech recognition to complex reasoning. To evaluate these models effectively, we should assess their ability to understand not only speech but also environmental sounds and paralinguistic features such as emotion, gender, and accent. Therefore, our evaluation involves designing benchmarks that cover three key aspects: 1) speech understanding, focusing on the interpretation of the semantic content within the audio; 2) audio scene understanding, focusing on non-human audio sounds; and 3) voice understanding, focusing on non-speech but human-related information and including paralinguistic tasks such as emotion, accent, and gender. For each task, we compile existing datasets and develop new ones tailored to the requirements for comprehensive understanding and reasoning. Figure 1 is an illustration of the tasks and datasets used in our setup.
@@
In AudioBench, the focus is on audio understanding of English semantics, where most models have not yet demonstrated satisfactory results. Meanwhile, multilingual understanding, which includes speech translation and code-switching, represents a useful application that could be integrated with future work.
@@
Category Dataset Name #Samples Hours Avg.L/Min.L/Max.L(s) Metrics Speech Understanding ASR LibriSpeech-Clean 2.6k 5.40 7.43 / 1.28 / 34.96 WER LibriSpeech-Other 2.9k 5.34 6.55 / 1.47 / 34.51 CommonVoice-15 16k 26.95 5.93 / 1.34 / 105.67 PeoplesSpeech 32k 59.20 6.54 / 1.00 / 99.91 GigaSpeech 18k 35.09 6.77 / 1.00 / 22.03 Tedlium3 1.1k 2.61 8.24 / 1.07 / 32.55 Tedlium3-Longform 11 2.61 0.8k / 0.3k / 1.6k Earning-21 44 39.26 3.2k / 1k / 5.7k Earning-22 125 119.88 3.4k / 0.87k / 7.4k SQA CN-College-Listen 2.2k 13.3 21.09 / 5.76 / 137.82 M.J. SLUE-P2-SQA5 408 4.5 39.85 / 13.00 / 40.0 DREAM-TTS 1.9k 18.1 34.14 / 3.14 / 261.85 Public-SG-SpeechQA 688 7.6 39.86 / 15.78 / 95.71 SI OpenHermes-Audio 100 0.16 5.95 / 2.04 / 15.77 M.J. ALPACA-Audio 100 0.12 4.32 / 1.80 / 8.85 Audio Scene Understanding AQA Clotho-AQA 2.2k 14.1 22.59 / 15.03 / 29.97 M.J. WavCaps-QA 304 0.87 10.28 / 1.0 / 30.62 AudioCaps-QA 313 0.86 9.86 / 3.27 / 10.00 AC WavCaps 1.7k 4.9 10.22 / 1.00 / 30.97 M.J. & AudioCaps 4.4k 12.1 9.86 / 1.74 / 10.0 METEOR Voice Understanding ER IEMOCAP-Emotion 1k 1.3 4.51 / 0.75 / 24.12 M.J. MELD-Sentiment 2.6k 2.4 3.35 / 0.13 / 304.9 MELD-Emotion 2.6k 2.4 3.35 / 0.13 / 304.9 AR VoxCeleb1-Accent 4.8k 11.2 8.27 / 3.96 / 69.04 M.J. GR VoxCeleb1-Gender 4.8k 11.2 8.27 / 3.96 / 69.04 M.J. IEMOCAP-Gender 1k 1.26 4.55 / 0.69 / 26.77
Table 1: Statistics of AudioBench Dataset Statistics. Yellow refers to our newly expanded and collected datasets to accommodate the missing of suitable datasets. WER refers to Word-Error-Rate. METEOR is a common metric for audio captioning. M.J. refers to model-as-judge, where we deploy Llama-3-70B-Instruct in the current setup.
@@ 3.2 Evaluation Setup
From a use case perspective, we expect LLMs to respond to user queries and generate natural responses. Therefore, we expect the model’s response to adhere to user queries and the generation style can adapt accordingly. Most conventional speech models and tasks have a constrained output space. Given classification tasks as an example, output is typically restricted to a finite set of categories or classes. However, the generation space of AudioLLMs is much broader and more complex, which could lead to difficulties in evaluation.
Therefore, we employed the Model-as-Judge (M.J.) approach for most tasks except for automatic speech recognition, where word error rate (WER) is used as the sole metric, and audio captioning, where the METEOR score is utilized as an additional measure. In subsequent experiments, we thoroughly validated the effectiveness of our proposed evaluation approach. We rescaled the M.J. scores to a 100-point scale for easier comparison.
Besides, we found that models can be less robust to diverse instruction, which can negatively impact their applicability and overall user experience. This issue is also present in pioneering text-based LLMs Zhu et al. (2023); Wang et al. (2024) but it is even more severe in AudioLLMs due to the complexity of processing multiple modalities during the fusion process. Therefore, we propose to leverage multiple instructions to evaluate a dataset, especially when the initial query lacks diversity. This approach is particularly relevant for tasks such as ASR, ER, AR, GR, and AC. The detailed analysis is presented in Section 4.
@@ 3.3.1 Speech Understanding
Dataset Name AudioLLMs Whisper+Llama3 SALMONN Qwen-Audio-Chat WavLLM Qwen2-Audio-Instruct Speech Understanding LibriSpeech-Clean 55.58 2.25 2.10 3.20 1.83 LibriSpeech-Other 41.80 4.16 4.80 6.07 3.71 CommonVoice-15 33.75 11.65 14.53 11.44 9.89 PeoplesSpeech 34.33 30.72 37.92 22.32 14.54 GigaSpeech 14.22 13.32 15.49 11.89 9.51 Tedlium3 8.56 4.00 6.62 6.39 3.81 Tedlium3-Longform 18.39 45.29 45.37 95.35 4.75 Earning-21 26.87 38.46 64.47 98.65 11.77 Earning-22 36.38 51.18 66.72 98.84 15.61 CN-College-Listen 50.51 60.85 65.43 74.50 85.25 SLUE-P2-SQA5 78.24 76.12 83.92 80.05 82.99 DREAM-TTS 55.93 57.76 64.56 66.70 86.09 Public-SG-SpeechQA 56.77 57.47 58.55 58.31 64.94 OpenHermes-Audio 19.20 11.00 22.40 44.80 63.0 ALPACA-Audio 12.40 9.60 21.60 52.60 70.8 Audio Scene Understanding Clotho-AQA 51.18 58.20 43.01 50.92 29.47 WavCaps-QA 46.25 38.68 26.25 44.47 17.38 AudioCaps-QA 47.03 47.99 29.84 45.75 16.71 WavCaps(M.J.) 21.16 29.25 6.40 33.78 3.45 AudioCaps(M.J.) 34.37 47.99 4.17 40.78 2.47 WavCaps(METEOR) 17.72 24.02 9.78 21.34 13.89 AudioCaps(METEOR) 21.20 27.70 6.70 19.89 7.95 Voice Understanding IEMOCAP-Emotion 21.56 27.34 45.91 49.30 34.43 MELD-Emotion 33.06 50.57 41.07 40.54 33.36 MELD-Sentiment 41.87 43.87 50.08 53.49 43.87 VoxCeleb1-Accent 28.06 45.70 37.65 29.19 39.33 VoxCeleb1-Gender 88.90 70.56 70.51 99.12 53.41 IEMOCAP-Gender 51.60 51.13 45.29 49.30 51.50
Table 2: Main results of four AudioLLMs and one cascade model.
Speech Question Answering (SQA). The task involves responding to questions based on speech-related audio content. However, there is a lack of suitable datasets. In this part, we involve both monologue understanding and dialogue understanding tasks. Three datasets are newly curated. Examples can be found in Table 5.
First, we collect questions from the English listening comprehension section of China’s national college entrance examinations, aimed at assessing students’ ability for listening comprehension in both academic and everyday situations. We manually compiled 271 questions and integrated them with 2000 questions sourced from Hu et al. (2024a), which includes questions from approximately 130 different exam sets. In line with the free-text QA format, we presented the questions without multiple-choice options. The dataset is named as CN-College-Listen.
The second dataset is DREAM-TTS, which is built upon text-based dialogue comprehension datasets DREAM Sun et al. (2019). We leverage one SOTA TTS engine Casanova et al. (2024) to convert the text input into the spoken format. We leverage 60 speakers in the generation and the genders are kept consistent with the dialogue content.
For Public-SG-SpeechQA dataset, we chose four public speaking videos from Singaporean politicians, accompanied by clean transcriptions. These transcriptions were manually segmented, and five questions per segment were generated using LLMs. Each question and its corresponding reference answer were manually reviewed to ensure their validity, where around 30% of the samples were discarded. Ultimately, this process resulted in the collection of 688 speech question-answer pairs for evaluation.
Speech Instruction (SI). This task evaluates whether the model can directly follow instructions provided through speech input, mirroring a natural human-computer interaction. Specifically, the question is delivered via audio, and the model is expected to understand and generate responses in the appropriate text format. This approach offers unique benefits. For instance, by incorporating paralinguistic information such as emotions, the model can adapt its responses based on the user’s emotional state. However, finding suitable testing examples has been a challenge. In this study, we synthesized audio instruction Casanova et al. (2024) from existing instruction-following datasets: ALPACA Taori et al. (2023) and OpenHermes Teknium (2023). We then involve humans in selecting instances to ensure that 1) the speech is accurately synthesized and 2) the content is suitable as spoken instructions. As a result, around 10.5% samples are kept and the examples can be found in Table 8.
Figure 2: Sensitivity tests were conducted on three different prompts on the same datasets. Performance is measured by WER (word error rate). The results state that SALMONN is the least robust when faced with varying prompt templates for the same task. Prompt 1-3: {"Turn the speech input into a text transcription", "Recognize the speech and give me the transcription", "Decode the audio and give me the written transcriptions."}